OpenStudy (anonymous):

Determine the vertex coordinates for: g(x) = (6 + x)(4 + x)

OpenStudy (anonymous):

= 24 +10x +x^2 vertex is at x = -10/2 = -5 coordinates are (-5,-1)

OpenStudy (anonymous):

thanks